OpenStudy (anonymous):

Simplify $\frac{x+1}{3}+\frac{2-3x}{2}$. Express your answer as a single fraction.

OpenStudy (mathstudent55):

Instead of $, use \( to start LaTeX.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Then use \) to end LaTeX instead of $.

OpenStudy (mathstudent55):

Also, I prefer dfrac for regular fractions and frac for fractions in exponents, numerators, and denominators. \(\dfrac{x+1}{3}+\dfrac{2-3x}{2}\)

OpenStudy (mathstudent55):

You are adding two fractions. The denominators are 3 and 2. What do you need to be able to add two fractions?

OpenStudy (anonymous):

make the denominators the same

OpenStudy (mathstudent55):

Correct. What is the LCD of 3 and 2?

OpenStudy (anonymous):

6

OpenStudy (mathstudent55):

Correct. The left fraction has a denominator of 3. We need 6, so we multiply it by 2/2. The right fraction has a denominator of 2. We need 6, so we multiply it by 3/3.

OpenStudy (mathstudent55):

\(\dfrac{2}{2} \cdot \dfrac{x+1}{3}+\dfrac{3}{3} \cdot \dfrac{2-3x}{2}\) Do you follow so far?

OpenStudy (anonymous):

I understand

OpenStudy (mathstudent55):

Now we multiply the numerators together and denominators together in both fractions. \(\dfrac{2(x+1)}{2 \times 3}+\dfrac{3(2-3x)}{3 \times2}\) We use the distributive property in the numerators to get: \(\dfrac{2x+2}{6}+\dfrac{6-9x}{6}\) Now we have a common denominator, so we just add the fractions by adding the numerators and setting that over the common denominator.

OpenStudy (mathstudent55):

\(\dfrac{2x+2+6-9x }{6}\) Now we need to combine like terms.

OpenStudy (mathstudent55):

\(\dfrac{-7x+8 }{6}\)